Stephen is 69 years younger than Wendy. 2 years ago, Wendy's age was 4 times Stephen's age. How old is Stephen now?
Found 2 solutions by VFBundy, greenestamps:
Answer by VFBundy(438)

Stephen's age now = s
Wendy's age now = s + 69

(s + 69) - 2 = 4(s - 2)

s + 67 = 4s - 8

-3s = -75

s = 25

Stephen's age now = s = 25

Answer by greenestamps(13200)



Of course there are many good ways to set up this problem for solving.

Following is what I would do....

Let x be Stephen's age 2 years ago; then Wendy's age then was 4x.

The difference in their ages 2 years ago was then 4x-x = 3x.

The difference in their ages (now, and 2 years ago, and always) is 69 years.

So 2 years ago the difference in their ages was 3x = 69.

So 2 years ago Stephen's age was x = 69/3 = 23 and Wendy's age was 4x = 92.

So now Stephen's age is 23+2 = 25 years.